OpenStudy (anonymous):

how does 4-(2sin(t))^2 get to be 4cos(t)^2

OpenStudy (anonymous):

\[\sin^{2} + \cos^{2} = 1 \] using this identity \[4 - (2 \sin(t))^{2} = 4 - 2^{2}(\sin^{2}(t))= 4-4(1-\cos^{2}(t)) = 4 - 4 +4\cos^{2}(t)\] \[4\cos^{2}(t)\]

OpenStudy (anonymous):

Thank you!